Journalist Report – December 13th

Crew 184 Journalist Report

Willie Schumann

13 December 2017

Title                            The Edge of Mars

Narrative                  Today for the first time, since our arrival, we could see a real dense set of clouds covering the Martian sky, which was really nice for a change. We jumped right into our space suits after breakfast and were poised to explore Matryoshka site six for more geological probes for earth. Yesterday our rover Deimos, named after our Martian Moon, was pretty worn out in the end of our EVA. So we decided to give him a little rest today and took its brothers Spirit and Opportunity out for a ride.

Both rovers are still very young and therefore un-experienced. We have to take them out from time to time for their batteries to grow stronger. Our exploration site today was the notorious Lith Canyon, which is very far north, basically at the edge of the Martian landscape that is still accessible for us considering our technical capacity. If we would go even further, we might not contain the power to come back in time to survive.

So it was a considerate risk to take the new rovers, but you have to stretch the range of the possible to progress. Space travel is not about always playing the save cards, it’s about to expand the borders of what mankind can achieve. And talking about the current limitations makes me refer to yesterdays evening. We opened a care package from earth with space food from Roskosmos, the Russian Space Agency. We got meat and cheese in tubes, which was quite alright, but there is definitely still room for culinary improvement. But as I am speaking Russian it was fun to get deep into the ingredients and share it with my fellow crewmembers.

We reached Lith Canyon pretty directly with no real detour. There are somewhat natural roads on Mars, shaped by wind and erosion. From experience we do not test our rovers to the extreme and keep them mainly on flat surfaces. That means, that we have to walk quite long distances through rough terrain. These longs walks on the other hand create other problems, but I will come to that in a little while.

The clouds over Lith canyon welcomed today’s EVA crew, consisting of First Officer Randazzo and Crew Engineer Hunt, in dramatic fashion and we found our designated sample sites very easily. The rocks, that were lined up on the walls of the Mountains looked like thin brittle plates, that were sprinkled over a desert. Almost like slate slabs made from very dense and compressed sand, very impressive.

But you have to remain very alert. When we walked closer to edge of the area we could look very deep into the maw of the canyon. You really have to watch your feet so you don’t accidently step on a sandy slab, which cracks and makes you slide. We always kept enough distance to the edge and backed one another up.

For Randazzo and Hunt it was actually the last EVA for a long time. They will be needed for other duties in the weeks to come. So there was a certain melancholy lingering in the air, complimented by the cover of clouds. We decided to shoot a remote control picture to commemorate the moment. And for the fact, that it was almost shot blindly from the ground it turned out brilliant.

So we finished our daily mission and headed back to the rovers. I was always a little bit behind, because the sun breaking through the net of clouds captivated me for additional shot. My sight was a bit blurred from my heavy breath, caused by the intensity of hiking up and down the canyon. I was lost for a brief moment, but reunited with the crew through radio communication.

But we all were disoriented and didn’t recall the position of our vehicles. Because of the rather long walk and the time we had spent in the canyon we couldn’t remember were we came from. So we chose the tallest crewmember around to go to a lookout to help us, who was obviously Crew Engineer Hunt. Did I mention, that we nicknamed him Big Foot? Now you know why!

Big Foot proved his value once again and got visual of our rovers. The sun was standing already low and we really had to hit the pedal to get to the hab before nightfall. Our young rovers contained their energy at a surprising high level. So it was quite a steep learning curve for the team and our vehicles, and the risk to take them out was proven justified. Another day on Mars has almost passed and we are all a bit wiser than before. Thank you for that dear red planet.

Personal Logbook             It was a great day on Mars with new visual impressions and experiences. The days now become shorter for us and the nights longer. As I am dependant on light I can not work us much as I want. I have to use my remaining time wisely to receive the results, that I want to achieve.

The group definitely grows closer together and knows one another inside out. I wonder how the dynamic of the entity that is crew 184 will develop in the future. For the time being we are happy for the time, that we can spend together.

Thank you very much for your help and attention.

Willie Schumann, Journalist, Crew 184

 

Sol Summary – December 13th

Summary Title:  In the final stretch (and clouds)

Mission Status:

We are getting close to the end of our mission, and the focus of the crew is maximizing the time we have left to complete our science objectives.

Today in the Matryoshka EVA series. Crewmembers Trisha Randazzo, Joshua Hunt, and Willie Schumann traveled to “Matrysoshka Site 6” which was located in the far north at the begging of Lith Canyon. After a long drive to the site with

We attempted to do solar observation tonight but were unfortunately stymied by extensive cloud cover all day long today.

This evening our team will focus on Anesthesia Experiment runs.  Once complete our team is currently hoping the clouds clear to allow us to watch the Geminid meteor shower which peaks tonight.  Shannon also allowed us to assemble her old telescope for use tonight.  Fingers crossed the clouds clear!

We preparing for our last ‘science day’ on Thursday, as we finish up both Matryoshka, Bricks, and Anesthesia experiments.  Friday is being held in reserve and is also allocated to cleaning the HAB and writing up lessons learned for the next crew.  The end is fast approaching!

Thanks for your help!

Sol Activity Summary:

  1. Matryoshka EVA Site #6
  2. ‘Bricks’ communication experiment
  3. Exercise
  4. General HAB Cleaning / Maintenance / Reports
  5. Telescope Assembly
  6. Anesthesia Experiment Runs this evening
  7. (Hopefully)  Geminid Observation!

Look Ahead Plan:

  1. Matryoshka EVA Site #7a
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. Anesthesia Experiment Runs
  5. Exercise
  6. General HAB Cleaning / Maintenance / Reports


Anomalies in work:

1.    HAB Leak from under EVA door.  No crew has showered yet today and a significant leak was observed by returning EVA crew.  It is likely this leak is coming from the Kitchen as the dishes were washed immediately prior to the crew returning.

2.  Power System Misconfiguration ongoing.

Weather: Mild and Cloudy

Crew Physical Status: All crew in good physical health.

EVA: See Above

Reports to be filed:

Greenhab

Operations

EVA Request

Sol Summary

Journalist

Astronomy

Support Requested:

1.  We have an extra CO detector that is marked ‘For Greenhab’.  The current detector in Greenhab is both a Smoke Alarm / CO Sensor.  Please advise whether you would like us to install a redundant CO Sensor in the Greenhab or return it to the HAB and mark it as a replacement part.

2.     Food Re-Supply

    1. 1 Box Raisin Bran
    2. 1 Box Powdered Milk
    3. 24 packs Instant Ramen
    4. 3 Boxes Pasta (any type)
    5. 2 Box Granola Bars
    6. 1 Box Saltine Cracker
    7. 1 Box Cracker (any type)
    8. 1 Box Hot Chocolate
    9. 1 Box Red Lobster Biscuits
    10. 1 Box Earl Grey Tea
    11. 1 Pack Chocolate Chips
    12. 1 Pack Tortillas
    13. 1 Pack Sour Cream\
    14. 1 Can SPAM
    15. 2 Can Chicken
    16. 1 Can Strawberries
    17. 1 Box Granola
    18. 1 Bottle Soy Sauce

Operations Report – December 13th

Crew 184 Operations Report 12/13/2017

SOL: SOL 12

Name of person filing report: JHunt

Non-nominal systems: None

Notes on non-nominal systems:  N/A

Generator (hours run): Generator turned off at 10:00am.

                                     Generator turned on at  5:10pm.

Solar— SOC @ 10:00 am : 100%

              SOC @ 5:10 pm: 68%

Diesel – 90%

Propane –  75 percent volume

Ethanol Free Gasoline (5 Gallon containers for ATV) – 0.5 Gallons

Water (trailer) – 400 gallons

Water (static) – 340 gallons

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 46 gallons

Toilet tank emptied: No

ATV’s Used: No ATVs used.

Oil Added? No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 00:00 hours

Notes on ATVs: ATVs were nominal

Deimos rover used: No

Hours: 00:00 hours

Beginning charge:

Ending charge:

Currently charging:Yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Yes

Hours: 01:00 Hours

Beginning charge: 100%

Ending charge: 55% 

Currently charging: Yes

Opportunity rover used:  Yes

Hours: 01:00 Hours

Beginning charge: 100%

Ending charge: 34%

Currently charging: Yes

Curiosity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No 

Summary of internet: Still 1.0 GB of Bandwidth

Summary of suits and radios:

EVA Suit #6 – Is missing one of the back black straps that secure the collar piece for the helmet.

Summary of Hab operations: All nominal

Summary of GreenHab operations:  See Greenhab Report

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: Crew is Healthy

Questions, concerns and requests to Mission Support:

GreenHab Report – December 13th

GreenHab Report

Trisha Randazzo

December 13, 2017

Environmental control: (Choose which is appropriate and explain further if needed)

Ambient (no heating or/cooling)

Ambient with window/door open

·         Inside temperature high 35 low 20

·         Inside humidity: 16%

Heating

·         Functioning nominally

Cooling

·         Not available

Both heating and cooling

·         Nominal

Shade cloth on/off

·         On

Average temperature: (N/A until new sensor is delivered)

·         N/A

Changes to crops:  Note all emergence of seedlings, death of seedlings, etc

·         All of the tomato plants are healthy! The ones that were doing poorly after thinning have healthy leaves, strong stems, and good green color. Yay for 100% healthy plants!

·       Some of the beans plants have drooping leaves. They are still strong in the stems but the leaves are noticeably drooping. The color is still a strong green.

Daily water usage for crops:

·         Some of the soil was extra dry, so will water twice a day about 12 hour apart.

Time(s) of watering for crops:

·         14:00

Morning research observations: (Currently not operational)

·         N/A

Changes to research plants:

·         N/A

Daily watering and amount of water used:

·          3.0 gallons

Aquaponics:  (Currently not operational)

·         N/A

Narrative:  Any other information you want to share

·         N/A

Support/supplies needed:

·         N/A

Astronomy Report – December 13th

Name:   Thomas Horn    Crew: 184
Date: 12/13/17

Sky Conditions: Cloud Covered

Wind Conditions: Light Winds

Observation Start Time: 11:30am

Observation End Time: 2:30pm

Summary:   Unfortunately observing was not able to be done today due to extensive cloud cover.  We are hoping the cloud clears enough to allow us to watch the Geminids tonight.  In the meantime we used the procedure Peter sent out yesterday to re-process the images from yesterday and it worked great, example below.

Thanks for the help!  We will try again tomorrow and hope the clouds clear.

Objects Viewed: Nothing



Problems Encountered: Clouds